OpenStudy (anonymous):

6^2x+4 = 3^3x . 2^x+8. Solve.... Plz

OpenStudy (anonymous):

6^2x+4 = (3^3x )( 2^x+8)

OpenStudy (anonymous):

\[6^{2x+4} = (3^3x)(2^{x+8})\]

OpenStudy (anonymous):

yup! n BTW this related with inequalities !

OpenStudy (anonymous):

help me fast plzz!

OpenStudy (radar):

(2x+4)log 6=(27x) 2^(x+8) (2x+4)log 6= log 27x + (x+8)log 2 Sorry, but I can't figure it out. Repost it and get some more help.

OpenStudy (radar):

@dumbcow, can you do this for dyanesh?

OpenStudy (anonymous):

Is it 3^2 * x or 3^(2x)?

OpenStudy (dumbcow):

@radar has the right idea, im assuming thats a typo and is should be 3^(3x) with x in the exponent if x is not an exponent then it can't be solved algebraically i will continue where radar left off \[(2x+4)\log 6 = x \log 27 + (x+8)\log 2\] \[2x \log 6 -x \log 27 -x \log 2 = 8 \log 2 -4 \log 6\] \[x (2 \log 6 -\log 27-\log 2) = 8 \log 2 -4 \log6\] \[x = \frac{8 \log2-4 \log6}{2 \log 6 -\log 27-\log 2} = 4\]

OpenStudy (radar):

Thanks @dumbcow, you made it look simple.